OpenStudy (anonymous):

Solve for x. 3x - 5x + 8x + x = 42 6 7 8

OpenStudy (anonymous):

Combine the x-terms together

OpenStudy (anonymous):

so ^6

OpenStudy (anonymous):

combine like terms.

OpenStudy (ivettef365):

add all the digits that are next to the variable, what do you get ?

OpenStudy (anonymous):

6

OpenStudy (anonymous):

?

OpenStudy (anonymous):

or is it 16

OpenStudy (anonymous):

3x - 5x + 8x + x = 42 x(3 - 5 + 8 + 1) = 42 7x = 42 x = 42 / 7 = 6

OpenStudy (anonymous):

so 6 ??????????????/

OpenStudy (anonymous):

x=6